# Yantao Linormal-$\diamondsuit$   Jianbing Zhangnormal-$\diamondsuit$   Zhiyong Wunormal-$\heartsuit$

### 17 Jan 2024 v1

@@ 5 Experiments
In this section, we first evaluate the GUI grounding capabilities of representative LVLMs and our proposed SeeClick on ScreenSpot. Subsequently, we adapt SeeClick to three downstream GUI agent tasks, exploring and analyzing the performance and potential of vision-based agents in various GUIs.
@@ 5.1 GUI Grounding on ScreenSpot
As the foundation of visual GUI agents, GUI grounding has not received extensive attention in current LVLMs evaluations (Liu et al., 2023b; Yu et al., 2023). Therefore, we compare various LVLMs on our specially constructed GUI grounding benchmark ScreenSpot, to examine their potential for GUI automation.
Compared LVLMs & Evaluation. We primarily evaluated two types of LVLMs: (1) versatile general LVLMs capable of tasks such as dialogue, recognition and grounding, including MiniGPT-v2 (Chen et al., 2023a) and Qwen-VL (Bai et al., 2023); (2) recent LVLMs specifically designed for GUI tasks, including Fuyu (Bavishi et al., 2023), CogAgent (Hong et al., 2023) and our SeeClick.
Considering that GUI agents require clicking on the correct position, we calculate the click accuracy as the metric, defined as the proportion of test samples where the model’ predicted location falls in the groundtruth element bounding box (Zhang et al., 2018; Li et al., 2022; Zhang et al., 2023b). We take the models’ generated point or the center of the generated bounding box as the predicted click location.
Results. As shown in Table 1, while general LVLMs have excelled in natural image grounding (e.g., RefCOCO (Kazemzadeh et al., 2014)), their GUI grounding performance on ScreenSpot is poor due to the significant differences between GUIs and natural images.
Accordingly, LVLMs trained specifically on GUI data have significant improvements. Our SeeClick achieved the best average performances across three GUI platforms and two types of elements, even with fewer model parameters and GUI-related training data than CogAgent. This demonstrates the efficiency of our GUI grounding pre-training; with the rich UI elements and diverse instructions collected from the web and mobile, SeeClick quickly learns to understand human instructions and locate elements, even in completely unseen scenarios like iOS and desktop. Notably, all models struggle with locating icons/widgets, highlighting the difficulty of identifying and grounding non-text elements on GUIs, and showcasing the unique challenges posed by ScreenSpot.
@@ 5.2 Visual GUI Agent Tasks
This section explores the capabilities of SeeClick when migrated to three downstream GUI agent tasks: MiniWob, AITW, and Mind2Web. Across these tasks, SeeClick emulates human behavior, determining the next action solely by observing interface screenshots, without needing any additional GUI information. On all three tasks, we trained SeeClick on the respective training splits and tested it on the test sets. Appendix B provides examples of SeeClick’s interactions in these tasks.
Figure 4: Comparison of SeeClick and LVLM baseline Qwen-VL on MiniWob. Tasks with yellow shadows indicate those with dynamic webpage layouts and element positions (closer to the real-world application of GUI agents, see more details in Section B.1). SeeClick outperformed Qwen-VL in most tasks, especially in those with dynamic elements, highlighting the effectiveness of our GUI grounding pre-training.
@@ 5.2.1 MiniWob
MiniWob Shi et al. (2017) comprises about 100 types of web automation tasks, where the agent is asked to interact with a simplified web environment to accomplish human instructions. Each task can templatize random variants and corresponding instructions controlled by a random seed, creating up to billions of possible task instances.
Existing open-source training data often lacks corresponding interface screenshots (Furuta et al., 2023). Therefore, we rollout 50 successful episodes using an LLM strategy for each task in (Zheng et al., 2023), resulting in a 2.8K episodes dataset to train SeeClick.
Compared Methods & Evaluation. We compared SeeClick with a range of offline training methods. Among these, the state-of-the-art method WebGUM (Furuta et al., 2023) uses screenshots as auxiliary information but still selects HTML elements as actions. Pix2Act (Shaw et al., 2023) is the only prior vision-based approach, trained with extensive data to perform clicking and typing actions on MiniWob. To verify the effectiveness of GUI grounding pre-training, we also report the results of the LVLM baseline Qwen-VL when trained with the same 2.8K dataset.
Due to the variance in evaluation task sets among different methods (Liu et al., 2018; Furuta et al., 2023; Shaw et al., 2023), for fairness, we report performance in two groups based on the overlapping MiniWob tasks with their approaches. We compute the success rate over 50 random seeds for each task, and then compute the mean over all MiniWob tasks as the final score.
@@
SeeClick Image 2.8K 68.0%
Table 3: Average scores of different methods on AITW. ClickAcc calculates the accuracy of click operation. The best results in each column are bold. SeeClick exhibits the best performance among both API-based LLMs (top row) and trained LVLMs (bottom row).
@@
Qwen-VL Image 53.3 61.2 56.3 61.7 52.0 56.9 60.1
SeeClick Image 56.0 64.5 57.7 63.6 57.3 59.8 68.4
Results. As depicted in Table 2, our purely vision-based approach SeeClick surpassed a series of HTML-based and image-based baselines with only a fraction of the training data. Notably, with an equivalent amount of training data (2.8K), SeeClick outperformed the previous offline sota WebGUM, which uses both HTML and screenshots as input. In comparison with visual approaches, thanks to the powerful reasoning and planning abilities of the LVLM and the localization skill imparted by our GUI grounding pre-training, SeeClick exceeded the sota visual method Pix2Act, using less than $0.3\%$ training data.
Furthermore, SeeClick significantly surpassed the LVLM baseline Qwen-VL by 20 percentage points, verifying that enhancing GUI grounding capabilities plays a crucial role in improving LVLMs’ performance on MiniWob tasks. To analyze in detail, we provide task-level comparisons between SeeClick and Qwen-VL on 35 MiniWob tasks in Figure 4. SeeClick shows notable improvements over Qwen-VL in most tasks, particularly in those where interface layouts and element positions dynamically change. This confirms our hypothesis that general LVLMs struggle with accurately clicking on interfaces, and our GUI grounding pre-training significantly enhances this capability, leading to better performance.
@@ 5.2.2 AITW
We evaluate the effectiveness of SeeClick in smartphone environments with a recently proposed Android UI automation dataset Android In The Wild (AITW). It encompasses 30k distinct task instructions and corresponding 715k human-annotated operation episodes. Previous approaches (Rawles et al., 2023; Zhan and Zhang, 2023) split AITW episode-wise into train/val/test (80/10/10%). However, with an average of over 20 episodes per instruction in AITW, there is a significant risk of overfitting due to: (1) extensive repetition in training trajectories; (2) instructions in the test set have appeared during training. In this work, we propose to split the training and testing sets by instruction. Specifically, we selected 545/688/306/700/700 instructions from General/Install/GoogleApps/Single/WebShopping respectively and retained only one annotated trajectory for each instruction. Then for each scenario, we selected 80% for training and the remaining for testing. This setting better reflects the performance of the trained agent in the real-world application.
Compared Methods & Evaluation. We compare SeeClick with two types of baselines: (1) API-based LLMs employing in-context learning or Chain-of-Thought (CoT), such as ChatGPT-CoT (Zhan and Zhang, 2023), PaLM2-CoT (Rawles et al., 2023) and the latest GPT-4V (Yan et al., 2023); (2) our own trained LVLM baseline Qwen-VL. We follow Rawles et al. (2023) to adopt the screen-wise action matching score as the main metrics, and compute the click accuracy (ClickAcc), which calculates the accuracy when both the reference and prediction are click actions.
Results. As illustrated in Table 3, SeeClick achieved the best average performance among both API-based LLMs and trained LVLMs. Specifically, SeeClick exhibited an 8 percentage point improvement in click accuracy over the LVLM baseline Qwen-VL, substantiating the hypothesis that GUI grounding significantly augments the precision of click actions in visual GUI agents, thereby improving their overall task performance.
5.2.3 Mind2Web
To assess SeeClick’s capabilities in web environments, we utilize the recently introduced Mini2Web dataset (Deng et al., 2023). Mind2Web comprises over 2000 open-ended tasks collected from 137 real websites, each with a high-level instruction and a human action trajectory sequence involving clicking, selecting and typing. Mind2Web was originally designed to evaluate text-based web agents, focusing on identifying actionable elements by selecting them from simplified HTML documents. This work explores visual web agents that predict click positions directly from interface screenshots. For this purpose, we augmented Mind2Web’s trajectory data with interface screenshots and bounding boxes for action elements at each step, sourced from their raw dump data (Deng et al., 2023). To the best of our knowledge, this is the first attempt of web agents relying solely on screenshots as inputs for navigating real websites.
Table 4: Comparsion of different methods on Mind2Web. The best results in each column are bold. Improvements of SeeClick over the LVLM baseline are underline. The GUI grounding pre-training nearly doubled the step success rate of SeeClick on real websites.
@@ 5.2.3 Mind2Web
Qwen-VL 14.9 86.2 12.6 12.1 83.7 10.1 9.6 84.3 8.0
SeeClick 26.3 86.2 23.7 21.9 82.9 18.8 22.1 84.1 20.2
Compared Methods & Evaluation. We compare SeeClick with a range of text-based web agents (Deng et al., 2023) and our LVLM baseline Qwen-VL. Mind2Act employs a two-stage method, where a small LM first generates candidate elements from raw HTML, and then a large LM selects the target element using multi-choice QA. Mind2Act (gen) modifies the QA module to directly generate the target element instead. GPT-3.5 and GPT-4 adopt the same multiple-choice QA formulation as MindAct, and include three demonstration examples for in-context learning.
@@
Results. As displayed in Table 4, SeeClick nearly doubled the element selection accuracy and step success rate compared to the LVLM baseline Qwen-VL. This indicates that improvements in GUI grounding directly correlate with enhanced performance in downstream web agent tasks. Furthermore, text-based methods often have lower overall Op.F1 due to nearly 20% groundtruth elements being incorrectly filtered out in the candidate generation stage, which leads to a zero Op.F1.
Although SeeClick can operate without extra HTML information, its performance lags behind state-of-the-art HTML-based methods, as predicting click locations on web pages is much more difficult than choosing from HTML candidates. This underscores the challenge of grounding in complex interfaces, suggesting substantial room for improvements toward a visual GUI agent that is applicable in real-world scenarios.
@@ 6 Conclusion
Appendix A Examples of ScreenSpot
@@
Appendix B Case Study
B.1 MiniWob

### 23 Feb 2024 v2

@@ 5 Experiments
In this section, we first evaluate the GUI grounding capabilities of representative LVLMs and our proposed SeeClick. Subsequently, we adapt SeeClick to mobile and web agent tasks, analyzing the correlation between the advanced grounding capacity and downstream task performance, while exploring the potential of purely vision-based GUI agents.
@@ 5.1 GUI Grounding on ScreenSpot
As the foundation of visual GUI agents, GUI grounding has not received adequate attention in current LVLMs evaluations (Liu et al., 2023b; Yu et al., 2023). Therefore, we evaluate LVLMs on our GUI-specific benchmark ScreenSpot.
Compared LVLMs & Evaluation. We primarily evaluated two types of LVLMs: (1) Generalist LVLMs capable of tasks such as dialogue, recognition and grounding, including MiniGPT-v2 (Chen et al., 2023a), Qwen-VL (Bai et al., 2023) and GPT-4V; (2) Recently released LVLMs specifically designed for GUI tasks, including Fuyu (Bavishi et al., 2023) and CogAgent (Hong et al., 2023).
Considering that GUI agents require clicking on the correct position, we calculate the click accuracy as the metric, defined as the proportion of test samples where the model predicted location falls in the ground truth element bounding box (Li et al., 2022; Zhang et al., 2023). More details about evaluation on ScreenSpot is in Appendix B.
Results. As shown in Table 1, while generalist LVLMs have excelled in natural image grounding, their GUI grounding performance on ScreenSpot is poor due to the significant differences between GUIs and natural images. Even GPT-4V struggles with accurately locating screen elements.
In comparison, GUI-specific LVLMs have significant improvements. SeeClick achieved the best average performances across GUI platforms and two types of elements, even with fewer parameters than CogAgent. This demonstrates the efficiency of our GUI grounding pre-training; with the rich UI elements and diverse instructions collected from the web and mobile, SeeClick quickly learns to understand human instructions for element localization, even in completely unseen scenarios like iOS and desktop. SeeClick exhibits slightly inferior performance in locating text within desktop and web compared to CogAgent, possibly due to lower resolution and much smaller training data. Notably, all models struggle with locating icons/widgets, highlighting the difficulty of identifying and grounding non-text elements on GUIs, which is the unique challenge posed by ScreenSpot.
Figure 5: Comparison of SeeClick and Qwen-VL on MiniWob. Tasks marked with yellow shadows feature dynamic webpage layouts, simulating real-world GUI agent applications (details in appendix Figure 11). SeeClick outperformed Qwen-VL in most tasks, highlighting the effectiveness of GUI grounding pre-training.
@@ 5.2 Visual GUI Agent Tasks
This section explores SeeClick’s application to mobile and computer agent tasks: MiniWob, AITW, and Mind2Web. We trained SeeClick on the respective training splits and tested it on the test sets. Across these tasks, with provided instructions and memory of previous actions, SeeClick determines the next action solely by observing interface screenshots. The detailed task settings, action formats and interaction examples are in Appendix C.
@@ 5.2.1 MiniWob
MiniWob Shi et al. (2017) comprises about 100 types of web automation tasks, where the agent is asked to interact with a simplified web environment to accomplish human instructions. Existing open-source training data often lacks corresponding interface screenshots (Furuta et al., 2023). Therefore, we rollout 50 successful episodes using an LLM strategy for each task in (Zheng et al., 2023), resulting in a 2.8K episodes dataset to train SeeClick.
Compared Methods & Evaluation. We compared SeeClick with a range of offline training methods. Among these, the state-of-the-art method WebGUM (Furuta et al., 2023) uses screenshots as auxiliary input but still selects HTML elements as actions. Pix2Act (Shaw et al., 2023) is the only prior vision-based approach, trained with extensive demonstration data to perform actions. To verify the effectiveness of GUI grounding pre-training, we also report the results of the LVLM baseline Qwen-VL when trained with the same 2.8K dataset.
Due to the variance in evaluation task sets among different methods (Liu et al., 2018; Furuta et al., 2023; Shaw et al., 2023), for fairness, we report performance in two groups based on the overlapping MiniWob tasks. We compute the success rate over 50 random seeds for each task and then compute the mean over all tasks as the final score. We provided task-wise scores in Section C.2.
@@
SeeClick Image 2.8K 67.0%
@@
Qwen-VL Image 49.5 59.9 46.9 64.7 50.7 54.3 57.4
SeeClick Image 54.0 66.4 54.9 63.5 57.6 59.3 66.4
Table 3: Average scores of different methods on AITW. ClickAcc calculates the accuracy of click operation. The best results in each column are bold. SeeClick exhibits the best performance among competing baselines.
Results. As depicted in Table 2, purely vision-based SeeClick surpassed strong baselines with substantially less training data. Notably, with an equivalent amount of 2.8K training data, it outperformed the offline sota WebGUM, which uses both HTML and screenshots as input. Moreover, thanks to LVLM’s powerful reasoning and planning abilities and our GUI grounding pre-training, SeeClick exceeded the sota visual method Pix2Act, using less than $0.3\%$ training data.
Furthermore, SeeClick significantly surpassed the LVLM baseline Qwen-VL by nearly 20 percentage points, underscoring the importance of GUI grounding in boosting LVLM’s performance. To analyze in detail, we provide task-level comparisons in Figure 5. SeeClick notably excelled in tasks with dynamic interface layouts and element positions, confirming our hypothesis that general LVLMs struggle with accurately clicking, and SeeClick markedly improves this aspect.
@@ 5.2.2 AITW
We evaluate SeeClick in smartphone environments with Android automation dataset Android In The Wild (AITW) (Rawles et al., 2023), which encompasses 30k instructions and corresponding 715k operation trajectories. Previous approaches split train/val/test episode-wise, which poses a clear risk of overfitting due to: (1) instructions in the test set have appeared in training, and (2) an average of 20 similar trajectories per instruction. In this work, we opt for an instruction-wise split, with 545/688/306/700/700 instructions from General/Install/GoogleApps/Single/WebShopping respectively, and retain one trajectory per instruction. We selected 80% for training and the remaining for testing in each subset. This split avoids overfitting and reflects the performance of agents on unseen instructions. Further details and results on the original split are in Section C.3.
Compared Methods & Evaluation. We compare SeeClick with two types of baselines: (1) API-based LLMs such as ChatGPT-CoT (Zhan and Zhang, 2023), PaLM2-CoT (Rawles et al., 2023) and the latest GPT-4V (Yan et al., 2023); (2) Our trained LVLM baseline Qwen-VL.
We follow Rawles et al. (2023) to adopt the screen-wise action matching score as the main metric and additionally compute the click accuracy (ClickAcc), which calculates the accuracy when both reference and prediction are click operations.
@@
Qwen-VL 15.9 86.7 13.3 13.2 83.5 9.2 14.1 84.3 12.0
SeeClick 28.3 87.0 25.5 21.4 80.6 16.4 23.2 84.8 20.8
Table 4: Comparsion of methods on Mind2Web. The best results in each column are bold. Improvements of SeeClick over LVLM baseline are underline, with GUI grounding pre-training nearly doubling the step success rate.
Results. As illustrated in Table 3, SeeClick achieved the best average performance among both API-based LLMs and trained LVLMs. Specifically, SeeClick exhibited a 9% increase in click accuracy over Qwen-VL, supporting the idea that GUI grounding enhances agent task performance through precise clicking.
5.2.3 Mind2Web
To assess SeeClick’s capabilities in web navigation, we utilize the recently introduced Mini2Web dataset (Deng et al., 2023), which comprises over 2000 open-ended tasks collected from 137 real websites, each with high-level instruction and corresponding human action trajectory. Mind2Web was originally designed for text-based agents, which select actionable elements from simplified HTML in each step. This work explores visual web agents that predict click positions directly from screenshots. For this purpose, we parsed screenshots and target element bounding boxes from the raw dump of Mind2Web. To the best of our knowledge, this is the first attempt of web agents relying solely on screenshots as inputs for navigating real websites.
Compared Methods & Evaluation. We compare with html-based web agents Mind2Act (Deng et al., 2023) and our visual baseline Qwen-VL. Mind2Act employs a two-stage method, where a small LM first generates candidate elements from raw HTML, then a large LM selects the target via multi-choice QA; Mind2Act (gen) directly generates the target element instead. GPT-3.5 and GPT-4 adopt the same multiple-choice QA formulation and include three demonstrations for in-context learning.
@@ 5.2.3 Mind2Web
Results. As displayed in Table 4, SeeClick nearly doubled the Ele.Acc and Step SR compared to Qwen-VL. This indicates that SeeClick’s improvement in GUI grounding correlates with enhanced performance in web agent tasks. HTML-based methods yield lower Op.F1 as around 20% of groundturth elements are filtered out during candidate generation. Although SeeClick can operate without extra HTML information, its performance trails sota HTML-based methods, since predicting click coordinates is much more difficult than choosing from HTML candidates. This highlights the difficulty of grounding in intricate interfaces, suggesting substantial room for improvement in visual agents for real-world application.
5.2.4 Grounding and Agent Performance
To investigate the correlation between grounding and agent performance, we analyze the average score improvements of several SeeClick’s checkpoints on ScreenSpot and three downstream tasks. As depicted in Figure 6, enhanced GUI grounding capacity consistently boosts agent task performance, highlighting its crucial role in developing advanced visual GUI agents.
5.2.5 SeeClick as Unified GUI Agent
To access the potential of vision-based solutions in unifying GUI agent tasks, we evaluated jointly training SeeClick on three downstream tasks. As shown in Table 5, the unified model exhibited a slight performance decline, possibly due to the significant distinct interface of different GUIs.
Figure 6: The correlation between agent tasks performance improvement and enhanced grounding ability.
MiniWob AITW Mind2web
Qwen-VLseparate 48.4 54.3 11.5
SeeClickseparate 67.0 59.3 20.9
SeeClickunified 64.1 57.1 19.5
Table 5: Separate v.s. unified training performance.
@@ 6 Conclusion
Limitations
SeeClick currently simplifies the GUI action space to mainly focus on clicking and typing, excluding complex actions like dragging and double-clicking. Additionally, limited by the performance of open-source LVLMs, training on agent-specific data is necessary for SeeClick to execute multi-step tasks on interfaces like mobile and computer.
Ethical considerations
GUI agents are developed to automate tasks and enhance efficiency on digital devices. These technologies are especially significant for individuals with visual impairments. Here are some ethical considerations:
Privacy Issues. The operation of GUI agents involves accessing and interacting with user interfaces that may contain personal or sensitive information. Ensuring data protection and user consent are paramount to maintaining privacy integrity.
Safety in Read-World Interactions. When GUI agents interact with the real world, there’s a risk of unintended harmful actions. Ensuring these agents operate within safe parameters is crucial to prevent negative outcomes.
Bias. The development of GUI agents must address potential biases in their algorithms, which could result in unequal performance across different user groups or interface designs. Mitigating bias is essential for equitable access and effectiveness.
Addressing these concerns requires ongoing research and development efforts, ensuring that the benefits of GUI agents are realized without compromising ethical standards.
Appendix A Details of SeeClick Pre-training
A.1 Pre-training Tasks
SeeClick employs pre-training tasks as outlined in Table 6. For the grounding task, we incorporate two forms: predicting center point coordinates (text_2_point) and predicting bounding box (text_2_bbox). For the task of generating text for elements (similar to OCR), we also include two categories: predicting text based on center point (point_2_text, widget captioning) coordinates and based on bounding boxes (bbox_2_text). Our preliminary experiments indicated that predicting points was slightly better than bounding boxes, likely due to the variable sizes of UI elements. Consequently, we increased the proportion of data with point localization. Finally, about 1 million samples are used for the continual pre-training of SeeClick.
For tasks involving coordinates, positions are represented as either the point (x,y) or the bounding box of (left, top, right, down), where each value is a two-decimal place number in the range [0,1] indicating the ratio of the corresponding position to the width or height of the image. Figure 7 provides some examples of the pre-training data.
Domain Task Sample Num
Web text_2_point 271K
text_2_bbox 54K
point_2_text 54K
bbox_2_text 54K
Mobile text_2_point 274K
text_2_bbox 56K
UI summarization 48K
widget captioning 42K
General LLaVA 145K
Total 1M
Table 6: All training data used by SeeClick.
Figure 7: Examples of SeeClick pre-training tasks.
A.2 Training Configurations
We employed the aforementioned data for continual pre-training of Qwen-VL-Chat to develop SeeClick. To enhance LVLM’s understanding of GUI images, we unlocked the gradients of its visual encoder and applied LoRA for fine-tuning. We adopt AdamW as the optimizer and use a cosine annealing scheduler with an init learning rate of 3e-5 and a global batch size of 64. All training takes around 24 hours on 8 NVIDIA A100 GPUs.
Figure 8: SeeClick on ScreenSpot. Blue dashed boxes represent the ground truth bounding boxes, while green and red pointers indicate correct and incorrect predictions.
Appendix B ScreenSpot Annotation & Evaluation
B.1 Human Annotation
We convened four experienced annotators, all either Ph.D. or master students in computer science, proficient in using mobile phones and computers and familiar with GUI operations. Initially, we assigned different GUI types to the annotators, such as iOS, Windows, and Web. Then, annotators were required to capture screenshots during their routine use (e.g., various apps) and subsequently annotate the clickable regions of frequently interacted elements using bounding boxes with annotation tool  [^2] http://makesense.bimant.com. Finally, these annotators were instructed to write corresponding English text commands for the annotated screen elements. All annotated interfaces and operational elements were in English and post-processed to remove personal information.
B.2 Sample Showcase
@@
B.3 Evaluation Detail
For comparing baselines, we tested the models’ grounding capabilities using their officially recommended approach. For instance, with CogAgent, we randomly selected prompts from the official set provided, such as "What steps do I need to take to <instruction>? (with grounding)", then the output coordinates (or the centers of bounding boxes) were taken as predicted points. For GPT-4V, we follow Yang et al. (2023b) to enable it to locate screen elements based on instructions. SeeClick’s predictions with points were marginally better than bounding boxes, thus we selected point prediction for final evaluation.
B.4 SeeClick Case Study & Error Analysis
Figure 8 presents some examples of SeeClick on ScreenSpot. SeeClick can comprehend human instructions and accurately locate screen elements. To conduct a detailed analysis of localization performance, we quantified the distances between predicted points and ground truth (the center of target elements) in Figure 9. It’s noteworthy that even incorrect predictions mostly occur near the target bounding box, suggesting the model recognizes the target but needs improvement in fine-grained localization.
Figure 9: Distance distribution of prediction point to ground truth. Most incorrect predictions are also close to the answer, suggesting the model recognizes the target but needs improvement in fine-grained localization.
Appendix C Downstream Agent Tasks
In this section, we first detail the formulation of SeeClick as a visual GUI agent, then separately introduce the settings for three downstream tasks, and finally show SeeClick’s interaction cases with the GUI across these tasks.
C.1 Formulation of SeeClick as Visual GUI Agent
Action Space SeeClick involves common human-UI interaction operations. Following AITW, we assigned an action_type id to each action type for model prediction.
  • click(x,y): 4. A click action at (x,y), where each value is a [0,1] number indicating the ratio of the corresponding position to the width or height of the image.
  • type("typed_text"): 3. An action of typing a piece of text.
  • select("value"): 2. An action for selecting an option from a dropdown menu on a webpage.
  • swipe(direction): Swipe actions for the screen, swipe up/down/left/right are assigned the ids 1, 0, 8, and 9 respectively.
  • PRESS BACK: 5. The action for returning to the previous step.
  • PRESS HOME: 6. The action for returning to the homepage.
  • PRESS ENTER: 7. The action of pressing the ENTER key to submit input content.
The first two actions, clicking and typing, are universally applicable across various GUIs. The third action, select, is defined according to the specifications in Mind2Web. The latter four actions, along with two additional states, TASK COMPLETE and TASK IMPOSSIBLE, are defined following the AITW framework for Android environments.
Agent Formulation SeeClick is an autonomous agent capable of executing human instructions on GUIs. It takes as input the instruction, a screenshot of the current interface and a series of (k=4 in our setting) previous actions, to predict the next action to be taken. Specifically, SeeClick uses the following prompt to execute each step of the agent:
During training and testing, we organize the data by step into the format described above.
C.2 MiniWob
MiniWob is a classic simplified web agent environment, built on Chrome, allowing low-level operations such as clicking and typing. It comprises around 100 tasks, where each task can templatize random variants and corresponding instructions controlled by a random seed, creating up to billions of possible task instances. We use 50 successful trajectories for each task provided in (Zheng et al., 2023) for training and test each task with 50 random seeds, following standard practices.
We report the average success rate across random seeds and tasks, automatically provided by the MiniWob environment. A task is considered successfully completed if executed correctly, while incorrect executions or exceeding the maximum number of actions (set as 30 here) are counted as failures. For the baselines in Table 2, we use the task-wise scores provided in their papers to calculate the average score for tasks overlapping with SeeClick. We also provided a task-wise comparison in Table 8.
Gen. Inst. GApps. Sing. WShop. Ovr.
Auto-UI 68.2 76.9 71.4 84.6 70.3 74.3
CogAgent 65.4 78.9 75.0 93.5 71.1 76.9
SeeClick 67.6 79.6 75.9 84.6 73.1 76.2
Table 7: Comparison on the origin split of AITW.
CC-Net (SL) WebN-T5 WebGUM Pix2Act Qwen-VL SeeClick
Choose-date 0.12 0.00 0.13 0.06 0.0 0.02
Click-button 0.78 1.0 1.0 0.32 0.42 0.96
Click-button-sequence 0.47 1.0 1.0 1.0 0.08 0.86
Click-checkboxes 0.32 0.96 1.0 0.99 0.44 0.78
Click-checkboxes-large 0.0 0.22 0.99 1.0 0.0 0.02
Click-checkboxes-soft 0.04 0.54 0.98 0.91 0.06 0.22
Click-checkboxes-transfer 0.36 0.63 0.99 0.76 0.60 0.70
Click-collapsible-2 0.17 0.00 0.95 0.31 0.0 0.48
Click-collapsible 0.81 0.00 0.98 0.80 1.0 1.0
Click-color 0.82 0.27 0.34 0.88 0.96 1.0
Click-dialog 0.95 1.0 1.0 0.12 0.96 1.0
Click-dialog-2 0.88 0.24 0.43 0.73 0.84 1.0
Click-link 0.59 1.0 1.0 0.86 0.0 0.90
Click-option 0.21 0.37 1.0 0.0 0.70 1.0
Click-pie 0.15 0.51 0.99 0.81 0.16 0.80
Click-shades 0.04 0.0 0.0 0.76 0.0 0.02
Click-shape 0.11 0.53 0.72 0.19 0.04 0.52
Click-tab 0.95 0.74 1.0 0.54 1.0 1.0
Click-tab-2 0.27 0.18 0.95 0.52 0.0 0.60
Click-tab-2-hard 0.19 0.12 0.95 0.0 0.16 0.42
Click-test 1.0 1.0 1.0 1.0 1.0 1.0
Click-test-2 0.95 1.0 1.0 1.0 0.72 0.94
Click-widget 0.56 1.0 1.0 0.87 0.38 0.58
Count-shape 0.21 0.41 0.68 0.0 0.20 0.28
Copy-paste 0.04 - - - 0.96 0.80
Copy-paste-2 0.01 - - - 0.96 0.80
Email-inbox 0.09 0.38 0.99 - 0.08 0.80
Email-inbox-forward-nl 0.0 0.6 1.0 - 0.24 0.74
Email-inbox-forward-nl-turk 0.0 0.33 1.0 - 0.16 0.56
Email-inbox-nl-turk 0.05 0.23 0.98 - 0.40 0.68
Enter-date 0.02 0.0 1.0 0.59 1.0 1.0
Enter-password 0.02 0.97 1.0 - 1.0 1.0
Enter-text 0.35 0.89 1.0 - 1.0 1.0
Enter-text-dynamic 0.39 0.98 1.0 - 0.96 1.0
Focus-text 0.99 1.0 1.0 - 1.0 1.0
Focus-text-2 0.96 1.0 1.0 - 0.84 0.96
Find-word 0.05 - - - 1.0 0.10
Grid-coordinate 0.66 0.49 1.0 0.97 0.96 0.52
Guess-number 0.21 0.0 0.11 - 1.0 1.0
Login-user 0.0 0.82 1.0 - 1.0 1.0
Login-user-popup 0.02 0.72 0.99 - 0.86 0.98
Multi-layouts 0.00 0.83 1.0 - 0.44 0.72
Multi-orderings 0.0 0.88 1.0 - 0.42 0.86
Identify-shape 0.68 - - 0.94 1.0 0.68
Navigate-tree 0.32 0.91 1.0 0.07 0.60 0.82
Search-engine 0.15 0.34 0.96 - 0.56 0.84
Simple-algebra 0.03 - - 0.99 0.48 0.38
Simple-arithmetic 0.38 - - 0.67 0.92 0.78
Text-transform 0.19 - - 0.91 0.36 0.46
Tic-tac-toe 0.32 0.48 0.56 0.76 0.30 0.58
Unicode-test 0.86 0.64 0.54 0.98
Use-autocomplete 0.07 0.22 0.98 0.95 0.72 0.82
Use-slider 0.18 - - 0.69 0.38 0.32
Use-spinner 0.47 0.07 0.11 - 0.24 0.16
Read-table 0.01 - - - 0.90 0.72
Average 0.336 (55) 0.552 (45) 0.861 (45) 0.646 (35) 0.564 (55) 0.712 (55)
Table 8: Mean scores across 55 MiniWob tasks.
C.3 AITW
AITW is a recently collected dataset for Android smartphone automation, where each sample comprises an instruction and an action trajectory with screenshots. AITW is divided into five subsets: General, Install, GoogleApps, Single, and WebShopping, totally including over 30K instructions and 700K episodes.
Despite AITW’s large scale, as stated in Section 5.2.2, the current train-test split poses a significant risk of overfitting, leading to experimental results that do not accurately reflect an agent’s generalization ability in the real world. We also conducted experiments on SeeClick using the origin split, as shown in Table 7, SeeClick is comparable to CogAgent’s performance. We believe that due to the severe overfitting, designing new agent frameworks or enlarging model size is unlikely to yield much improvements on this split.
To address the aforementioned issue, we propose to divide the train/val/test in an instruction-wise manner. Specifically, we selected 545/688/306/700/700 instructions from the General/Install/GoogleApps/Single/WebShopping subsets, and retained only one annotated episode for each instruction. To avoid imbalance in joint training, we randomly chose 700 instructions from Single and WebShopping. Given the similarity among instructions within Single and WebShopping, these 700 instructions are representative of performance on these two subsets. Next, we allocate 80% for training and the remaining 20% for testing, and select additional 5*100 episodes to form the validation set from the origin data. The data used for training, validation, and testing will be open-sourced to serve as an effective evaluation.
The other settings are consistent with previous work, calculating a screen-wise matching score that considers both the correctness of the action type and its value (e.g., the click point or typed text). The screen-wise matching score is correlates with the task completion score judged by humans (Rawles et al., 2023).
C.4 Mind2web
Mind2Web is a recently proposed dataset for developing generalist web agents for real-world websites, originally designed for text-based agents. Therefore, the origin observation in each step only includes the HTML code of the current webpage. To train and evaluate visual-based agents, we extracted web screenshots and the bounding boxes of target operational elements for each step from Mind2Web’s raw dump. One issue with Mind2Web’s original HTML observation is that it captures the entire page, including scrolling, with its screenshots being long captures (e.g., 1920*12000). Predicting operational positions from such high-resolution long screenshots is impractical for current LVLMs and does not align with human operations. To address this, for target elements not at the top, we randomly crop around their location, maintaining a consistent screenshot resolution of 1920*1080 for all observed interfaces.
Mind2Web first calculates Element Accuracy (Ele.Acc) which compares the predicted element with groundtruth, and Operation F1 (Op.F1) which calculates the token-level F1 score for the predicted operation. Operation F1 is equivalent to the accuracy of click operations but takes into account the correctness of input values for type and select operations. For our vision-based approach, Element Accuracy is computed as the accuracy of predicted click points falling in the groundtruth elements’ bounding box. Then, a step is considered successful (Step SR) if both the predicted element and operation are correct.
C.5 Case Study